OpenStudy (anonymous):

Christina has some money at home, and the amount is modeled by the function h(x) = 103. She read about a bank that has savings accounts that accrue interest according to the function s(x) = (1.03)x − 1. Using complete sentences, explain to Christina how she can combine her functions to create a new function, and explain what this new function means.

OpenStudy (anonymous):

@mathmate @mathrulezz

OpenStudy (anonymous):

@amorfide last one, don't worry if you don't know it... Ill give you a medal anyway! Sorry

OpenStudy (amorfide):

has money h(x)=103 saving account s(x)=1.03x-1 already has h(x) as she starts with money at home (103) she wants to use a savings account to gain interest s(x) she can combine the function by depositing her money into an account, and gaining interest this would mean depositing the function h(x) into s(x) which would give s(h(x)) s(h(x))=1.03(h(x))-1 s(h(x))=1.03(103)-1

OpenStudy (amorfide):

this new function would give the amount of money in the savings accoutn after gaining interest

OpenStudy (anonymous):

Thank you